OpenStudy (anonymous):

the average of 15 numbers is 7. Angelo is adding these numbers and mistakenly reads 6 of the numbers as 9 instead of 7,what average will he get??..help plz..

OpenStudy (dumbcow):

this means six times he adds 2 more than he should....total of 12 so the sum is 12 higher than it should sum = 15*7 - 12 = 93 avg = 93/15 = 6.2

OpenStudy (anonymous):

6 times 2??why??

OpenStudy (kropot72):

\[average=\frac{93}{15}=6\frac{3}{15}=6.2\]